OpenStudy (anonymous):

What is the Least Common Denominator of the fraction 7 over 3 x plus the fraction 2 over x plus the fraction x over x minus 1 ?

OpenStudy (anonymous):

Is this the question? \[\frac{7}{3x} + \frac{2}{x}+\frac{x}{x-1}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The denominators have no common factors. The must be multiplied together.

OpenStudy (anonymous):

3x - 1 3(x2 - 1) 3(x - 1) 3x(x - 1) these are the choices they gave me

OpenStudy (anonymous):

I see it now. The \(3x\) and \(x\) have a common denominator or \(x\) therefore, you only need to multiply the first and third denominators.

OpenStudy (radar):

I like the last one, let me check further.

OpenStudy (radar):



OpenStudy (anonymous):

So it is the last one?

OpenStudy (anonymous):

Yes, it is. Do you understand why?

OpenStudy (anonymous):

es i do thank you guys :)

OpenStudy (anonymous):

Yes*

OpenStudy (anonymous):

Great!

OpenStudy (radar):